9. Bakugo is initially at rest but then darts forward 800 min 125. What is his acceleration? Formula: x= 1/2 at2


8. A Ford Mustang racing at NASCAR accelerates at a rate of 60 m/s for 2.0s. How far does it move Formula: x= 1/2 at2


7.. An object initially at rest experiences an acceleration of 9.8 ms. How much time will it take to achieve a velocity of 49 m/s?

Formula: t = v - v0

-------

a



1
Expert's answer
2022-01-14T09:45:51-0500

9. "a=2x/t^2=2·800/125^2=0.102\text{ m/s}^2."

8. "x=at^2/2=60·2^2/2=120\text{ m}."

7. "t=(v-v_0)/g=(49-0)/9.8=5\text{ s}."